# DIRAS: Efficient LLM-Assisted Annotation of Document Relevance in Retrieval Augmented Generation

### v1

@@ Appendix I ClimRetrieve Dataset Overview
While the question itself only addresses "climate change adaptation", the relevance definition allows for contents that are in line with "climate change mitigation". Mitigation is much broader and much more discussed in sustainability reports. The examples in the definition additionally broaden the scope. This also hints at the third avenue to explain the many false negatives: they contain a lot of partially relevant (document, query) pairs. This can also explain the significant drop in average confidence between true and false negatives in Table 14. However, the important aspect of the DIRAS pipeline is that it is consistent with the provided definitions which Fig. 16 indicates.
This investigation unveils clear edge cases for partial relevance and the mismatch of different definitions, i.e. mental models of annotators. Nonetheless, Fig. 16 reveals that, even amongst these hard examples, the approach remains consistent with the provided definitions. Although looking at edge cases, our reannotations are largely aligned with our model when overturning the non-golden irrelevant documents of ClimRetrieve.
To further get an intuition on edge cases, we turn our attention to the false positives. These represent the most complicated cases where the ClimRetrieve human annotator deemed the (document, query) pair relevant and the model arose with the judgment of not relevant. Since only 148 (document, query) pairs are in the false positives, we randomly sample 96 (document, query) pairs and reannotate them. The first two quantitative observations remain similar to the false negatives. First, Table 14 shows that the average confidence for the false positives is significantly lower than those of the true positives indicating they present edge cases. Indeed, they have the lowest average confidence among all categories with 0.8871. However, Fig. 17 again shows that with a rising confidence threshold, the agreement between the prediction and our reannotation rises. This time, the level of partial relevance is higher than for the true negatives. These two cues towards higher complexity continue in the qualitative assessment of our reannotations. We repeat to ask the question why is there a discrepancy between our model and the ClimRetrieve annotations? This time, we identify a multitude of edge cases. First, a technical edge case comes into play. ClimeRetrieve was created by matching span-labeled relevant texts with documents retrieved from the reports. A document is deemed relevant when one sentence of the span-labeled text appears in a document. However, one sentence alone cut off or in a different context may not be relevant. In 19% (18 of 96) of the annotated samples, this is the reason for a wrong label assigned. Second, the nature of partially relevant labels seems to be ambiguous. While the ClimRetrieve human annotator deemed many samples partially relevant, our reannotation overturned some of these cases in line with our model. Nonetheless, there remains a significant chunk of partially relevant samples (see Fig. 17). Third, using our model, we overturn some of the decisions of the ClimRetrieve annotators entirely. Even if they deemed it relevant, our reannotation and model suggest irrelevance given the definition. This can again be linked to the mismatch of the exact mental model of the ClimRetrieve annotators and our approximated reannotation guidelines. It has to be noted that these decisions may again be overruled by a second or third reannotator. The samples generally all have a connection with the topics in the question. Rather than presenting a gold standard, this comparison allows us to understand the fine-grained nuances of relevance, therefore reinforcing the very need of this project. Clear misannotations were only marked in 3 of 96 cases where our annotator could not identify any remote relevance.
Finally, we investigate the large chunk of the (document, query) pairs, the true negatives (model says not relevant, human says not relevant). Having a large basis of 37.409 (document, query) pairs, we again sample 6 pairs per question using two samples with a Llama3-Ask confidence above 0.9, two between 0.9 and 0.7, and two below 0.7. The true negatives are not at all comparable with the investigations before because we now sample from pairs with apriori agreement. Thus, we don’t investigate edge cases in this category.
This is also confirmed by the results (see Fig. 18). Only two samples with a Llama3-Tok confidence threshold below 0.5 were labeled as partially relevant, none were labeled as relevant. This reinforces the general results in § 3 and shows that the previous two reannotations concerned edge cases.
Collectively, these observations prove the motivation for this project in developing a nuanced stand beyond binary and even partial relevance. They show that the models function consistently with themselves - even in edge-case scenarios. This can mean something different than consistent with human annotators who cannot share their mental model.
"Golden" model with
"golden" definition
True False
Human with own mental model True
447
148
False
5.441
37.409
Table 13: Comparison of the Llama-3’s relevance prediction vs. the ClimRetrieve human annotators on ClimRetrieve’s 43K (document, query) pairs.
Labels assigned by
fine-tuned LLama3
Average Confidence
Llama3-Tok
all 0.9681
positives 0.9049
negatives 0.978
true positives 0.9575
false positives 0.8871
true negatives 0.9784
false negatives 0.9006
Table 14: Average confidence scores of Llama3-Tok for the different classification categories.
relevance definition:
Meaning of the question: The question "Has the company identified any synergies between its climate change adaptation goals and other business goals?" is asking for information about how the companys efforts to adapt to climate change are aligned with, or can complement, its other business objectives. This includes understanding if climate change initiatives also support broader strategic goals such as cost reduction, risk management, innovation, market expansion, or reputation enhancement.
Examples of information that the question is looking for:
1. Descriptions of [...]
MISSED OUT DOCUMENT:
[...] Along with these efforts, in each business segment, Sony develops and enhances risk management and business continuity plans (BCPs) from the perspective of improving risk management across supply chains, through the identification, analysis, and assessment of business continuity risks. Flood damage has grown in recent years due to the impact of climate change, prompting Sony to reassess the flood risk at its manufacturing sites in Japan and implement preventative measures that will mitigate flood damage and facilitate rapid recovery. Sony is collaborating with relevant companies and organizations, and conducts hands-on drills to address foreseeable risks, in an effort to enhance business continuity and accelerate flood recovery. Sony will continue to increase its resilience to climate change, based on its analyses and initiatives.* A global initiative in which participating corporations aim to operate on 100% renewable electricity. It is headed by an international non-governmental organization, the Climate Group, in partnership with the CDP.
Figure 15: False Negative example of clearly relevant but missed out information in ClimRetrieve.
Figure 16: Percentage of agreement between model and our human annotation as well as the data remaining when adjusting the confidence threshold for false negatives.
Figure 17: Percentage of agreement between model and our human annotation as well as the data remaining when adjusting the confidence threshold for false positives.
Figure 18: Percentage of agreement between model and our human annotation as well as the data remaining when adjusting the confidence threshold for true negatives.

### v2

@@ Appendix I Creation of the Improved Relevance Definitions
DOCUMENT: Enabling a more sustainable world Microsofts actions alone will not solve the climate crisis. As a global technology leader, we are also committed to helping build the enabling societal conditions that will support a net zero economy. Were focused on accelerating the availability of new climate technologies, strengthening our climate policy agenda, helping to develop a more reliable and interoperable carbon accounting system, advocating for skilling programs to expand the green workforce, and working to enable a just energy transition.
Figure 15: Example for which DIRAS Llama3-Tok assigns "not relevant" and ClimRetrieve assigned "relevant". This example shows that while DIRAS Llama3-Tok might differ with the opinion of ClimRetrieve’s annotator, it is consistent with its own definition.
Appendix M Implementation Details of Experiments on QA Datasets
ALCE Data: We obtain ELI5, ASQA, and QAMPARI from ALCE (Gao et al., 2023), where they parse the original open-domain QAs into RAG forms [^11] Data files in https://github.com/princeton-nlp/ALCE. For each question, ALCE annotates 5 documents as oracle based on retrieval recall and reference answers, which are used as context relevance labels in our experiment. For each dataset, we randomly sample 100 queries to construct DIRAS training data, following the process in Fig. 2. Top-5 is selected for balanced sampling, thus resulting in 1000 (query, document) pairs for training. We sample 50 questions for test data, and include all (query, document) pairs for them, resulting in 5K (query, document) pairs for each dataset.
RAG-Bench Data: RAG-Bench classifies RAG sources into four types: (A) relevant and with answers, (B) relevant topic but without answers, (C) irrelevant topic, and (D) with counterfactual answers. We find (B) addresses partial relevance that DIRAS cares about. Therefore, we leverage its dev set for training and test set for testing, where (A) and (D) become relevant documents, and (B) and (C) are used as irrelevant ones.
Disagreement Sampling for Table 8: We sample 200 disagreed annotations four each dataset, 50 samples from each confidence range: Conf$<90$, $90<$Conf$<95$, $95<$Conf$<98$, and $98<$Conf$<100$ to balancedly cover different confidence scores (these bins are of similar size).
You will be provided with a <question> and a <paragraph>. Your job is to assess whether the <paragraph> is useful in answering the <question>, given the <background_information> defining what is useful.
<background_information>: "The <paragraph> is useful only if some of its content directly answer the <question> or at least a part of the <question>. Content with relevant topic but without direct answers are not useful."
<question>: "{question}"
<paragraph>: "{paragraph_chunk}"
Is <paragraph> useful for answering <question>? Provide your best guess and your confidence that the guess is correct. Reply in the following format:
[Guess]: <Your most likely guess, should be one of "Yes" or "No".>
Figure 16: The DIRAS prompt for QA experiments.